OpenStudy (anonymous):

Bill has brown eyes (Bb) and unattached earlobes (EE). He marries Henrietta who has green eyes (bb) and unattached earlobes(Ee). What is the probable phenotypic ratio of their offspring? What is the probable genotypic ratio of their offspring?

OpenStudy (jewlzme17):

BBEE= 1/16 BEEb= 2/16 EEbb=1/16 BBEe=2/16 BEbe= 4/16 Ebbe=2/16 BBee=1/16 Bbee=2/16 bbee= 1/16 Since both Earlobes They will come out the same For Brown eyes (Bb) and unattached earlobes (EE) = 9/16 (Ee)= 16/16 ( because it can be either dominant or recessive but only for BbEe) For Green eyes(bb) and unattached earlobes (EE) = 3/16 (Ee)= 4/16 (one is all recessive) Hope this helped!

OpenStudy (blues):

Bill is BbEE and Henrietta is bbEe. Bill's possible combinations of gametes are BE, BE, bE and bE. Henrietta's possible combinations of gametes are bE, be, bE and be. So you set up your Punnett Square:  Count the genotyptic ratios, convert them to phenotypes - note that sometimes DIFFERENT genotypes will actually produce the SAME phenotype - and figure out your ratios.